Ticket #322 (closed defect: fixed)

Opened 5 years ago

Last modified 4 years ago

40G testing OSF-SL-UvA

Reported by: ibaldin Owned by: ckh
Priority: major Milestone:
Component: Infrastructure: Rack hardware Version: baseline
Keywords: Cc: vjo

Description

To keep track of 40G issues.

Latest status:

1. Waiting on decision which optics/active cabling to use in ALU-IBM link at OSF

2. Testing from SL to UvA - unknown.

Change History

Changed 5 years ago by ibaldin

  • keywords vjo added

Changed 5 years ago by ibaldin

  • cc vjo added
  • keywords vjo removed

Changed 5 years ago by ibaldin

OSF is ordering SR part.

Changed 4 years ago by ibaldin

  • status changed from new to closed
  • resolution set to fixed

OSF-SL have been tested to provide 29.9G (constrained by the speed of bare-metal hardware in the racks). The repeatable process is being established.

Note: See TracTickets for help on using tickets.